### v2

# The evolution of trust as a cognitive shortcut in repeated interactions

Abstract
Trust is often thought to increase cooperation. However, game-theoretic models often fail to distinguish between cooperative behaviour and trust. This makes it difficult to measure trust and determine its effect in different social dilemmas. We address this here by formalising trust as a cognitive shortcut in repeated games. This functions by avoiding verifying a partner’s actions once a threshold level of cooperativeness has been observed. We consider trust-based strategies that implement this heuristic, and systematically analyse their evolution across the space of two-player symmetric social dilemma games. We find that where it is costly to verify whether another agent’s actions were cooperative, as is the case in many real-world settings, then trust-based strategies can outcompete standard reciprocal strategies such as Tit-for-Tat in many social dilemmas. Moreover, the presence of trust increases the overall level of cooperation in the population, especially in cases where agents can make unintentional errors in their actions. This occurs even in the presence of strategies designed to build and then exploit trust. Overall, our results demonstrate the individual adaptive benefit to an agent of using a trust heuristic, and provide a formal theory for how trust can promote cooperation in different types of social interaction. We discuss the implications of this for interactions between humans and artificial intelligence agents.
Keywords: Trust modelling, Costly monitoring, Evolutionary game theory, Multi-agent systems, Social dilemma.
1 Introduction
Trust underlies cooperation in every human society [1, 2], and hence has been called “the glue of healthy societies and the grease of economic productivity” [3]. This implies that the Artificial Intelligence (AI) agents of the future will also need the capability to build trust, in order to cooperate effectively both with each other and with people [4, 5]. But this raises a fundamental question: what exactly is the individual benefit of trusting another agent? By trusting another, an agent exposes itself to risk [6] – the risk that the other agent might not act in an honest and reliable way.
This question has been difficult to answer because there are many different verbal definitions of trust in both social science and artificial intelligence. These definitions can be categorised into: 1. behavioural definitions, which equate trust with more cooperative choices in a game, and 2. psychological definitions, which equate trust with psychological states such as viewing another individual as benevolent and acting with integrity [7]. Prior work on trust in economics [8] and multi-agent systems [9] has typically followed 1., and effectively equated trust with high levels of cooperation. The risky investment decision modelled by the “trust game” [10, 11, 12], which is effectively a one-sided Prisoner’s Dilemma [13], provides a paradigm example. The fundamental problem with this is that it conflates an outcome, cooperation, with one of its causes, trust [14, 15, 16]. As such, these definitions of trust do not allow us to delineate the role of trust in promoting cooperation. Although another (smaller) tradition in multi-agent systems has followed 2., and equipped agents with cognitive mechanisms to make trust decisions [17], this has still not resulted in a general and mechanistic way to measure trust that applies in a wide range of settings. This is because it defines trust relative to an agent’s internal state, which may very well not be observable in an interpretable way, e.g. as is the case with current Large Language Model (LLM) agents, as well as for human agents.
To overcome the problems with both 1. and 2., we need a measure of trust that is behavioural, so that it can be observed regardless of the agent’s architecture (artificial or human), but that does not conflate trust with cooperation. To do so, in this paper we develop an objective measure derived from the adaptive function of trust behaviours, i.e. the individual benefits that trust brings to an agent. We do this by using evolutionary game theory (EGT) to formalise a key theory of trust from social science, originating from the work of Luhmann [6]. Luhmann’s theory defines trust as a heuristic – a decision-making shortcut that people use in social interactions – to avoid having to constantly reason about the likelihood of every possible outcome [6, 7]. When trust is established, the benefit is that people save the time and effort of trying to exhaustively calculate whether or not they should cooperate and take the risk of being exploited.
We turn Luhmann’s theory into a behavioural measure of trust that can apply to both human and artificial agents by building on the theory of repeated games. In repeated games, agents interact with each other several times (rounds). This gives agents the possibility to condition their behaviour on how the other agent has acted previously. For example, by using reciprocal strategies such as “cooperate this time if the other agent cooperated on the previous round”. These reciprocal strategies – where agents can base their knowledge of how the other agent has acted previously on either their own first-hand knowledge (direct reciprocity) or information from third parties (indirect reciprocity) – underpin cooperation in human societies [18]. But crucially, to be able to use reciprocal strategies, an agent needs to know whether the observed actions of another agent are actually cooperative or not. It is traditionally assumed that once an agent observes an action, it can instantly and with zero cost determine whether that observed action was cooperative. However, in reality, verifying whether an observed action was actually cooperative carries an opportunity cost, representing the time and effort spent. For example, in e-commerce it may not be easy for a buyer to tell whether a seller sent them goods of the advertised quality, rather than say a fake, and so choosing to verify this may take a substantial amount of their time. Similarly, verifying the actions of AI agents may involve simulating execution of their source code [19], the cost of which would depend on the power of the simulating agent and the auditing frameworks in place.
Classic game theory has traditionally studied this as a problem of costly monitoring. Some models have considered cases where agents can choose to pay a cost to reveal their co-players’ actions, and have demonstrated folk theorem results about the existence of cooperative equilibria under particular conditions, such as communication [20] or the presence of a random signal [21, 22]. However, these models have focussed on equilibrium outcomes rather than competition amongst different strategies employed by cognitively-bounded agents. Consequently, they have not connected costly monitoring with an individual trust heuristic. By making this connection, we argue that trust can be objectively measured by observing how often an agent chooses not to pay a verification cost, and hence does not invest time and effort in checking whether its fellow agents are cooperating.
We operationalise this functional definition of trust into an agent’s decision-making mechanism by introducing a class of trust-based strategies into repeated games. In contrast to traditional reciprocal strategies such as Tit-for-Tat [23], trust-based strategies only occasionally verify their partner’s actions once a trust threshold has been reached, i.e. after their co-player has made actions that have been verified to be cooperative for a certain number of rounds. By doing so, these strategies reduce the opportunity cost of verifying the action of their partner every round. We previously showed that trust-based strategies evolved in a repeated Prisoner’s Dilemma game, and that their presence increased the frequency of cooperative actions in the overall population of agents. This result held for a wide range of non-zero opportunity costs [24]. But this work left unclear the extent to which this explanation for the evolution of trust would generalise to other common types of social interaction beyond the repeated Prisoner’s Dilemma.
We resolve this here by examining the extent to which trust-based strategies evolve and promote cooperation in a wide range of social interactions, including the full space of symmetric two-player social dilemma games [25, 26], and in settings where agents can make unintentional mistakes in their actions. We find that a trust heuristic is most effective at promoting cooperation in situations where there is a strong temptation to defect. In contrast, trust has the least effect on promoting cooperation in coordination interactions (Stag-Hunt games), although it still evolves in these cases. But in all cases, we find that trust is beneficial in promoting cooperation if there is a moderate probability of agents making unintentional errors in their actions. Overall, our results show that a wide range of repeated social dilemmas select for the evolution of a trust heuristic that avoids always verifying another agent’s actions.
2 Models and methods
We first describe the model, and then present the methods used to analyse it. We follow the presentation in the Models and methods section of [24], but extend the model to cover different types of game, and introduce the possibility of agents making errors in their actions.
2.1 Pairwise social dilemmas
We model the interactions between individuals as a pairwise social dilemma, defined by the following parameterised payoff matrix (for row player), as in [25, 26]:
[…]
An agent who chooses the cooperate (C) action when their partner chooses the defect (D) action receives the sucker’s payoff $S$, while the agent that defects gains the temptation payoff, $T$. Mutual cooperation (resp., defection) gives a reward $R$ (resp., punishment P) for both agents. This payoff matrix defines the stage game, i.e. the payoffs that result from a single round of the interaction.
Depending on the ordering of the four payoffs, different social dilemmas arise [27, 26]. We use the following scaling of the payoff matrix, where $R=1$, $P=0$, $0\leq T\leq 2$, $1\leq S\leq 1$. Varying $S$ and $T$ capture all possible pairwise social dilemmas. Namely, (see [26])
  • 1.
    Prisoner’s Dilemma (PD): $T>1>0>S$;
  • 2.
    Snowdrift (SD): $T>1>S>0$ and $T+S<2$;
  • 3.
    Stag Hunt (SH): $1>T>0>S$.
In a repeated game, agents have strategies. These are functions which map from the information that the agent has about the actions its partner has made on previous rounds, to the action that the agent will take in the current round. In a population of $N$ individuals playing a repeated social dilemma as defined above, the fitness of an individual with a strategy A in a population with $k$ As and (Nk)(N-k) Bs is given by
$\Pi_{A}(k)=\frac{1}{r(N-1)}\sum_{j=1}^{r}[(k-1)\pi_{A,A}(j)+(N-k)\pi_{A,B}(j)],$ (1)
where $\pi_{A,A}(j)$ and $\pi_{A,B}(j)$ denote the payoff obtained by the A individual in a single round, $j$, when it interacts with an individual of strategy A (B, respectively), as defined by the payoff matrix above. The symbol $r$ denotes the total number of rounds of the game. Rather than considering a fixed number of rounds that agents know in advance, we instead make the standard assumption that the game continues for another round with probability $w$. This results in an expected number of $r=(1-w)^{-1}$ rounds per interaction [28]. Throughout this paper, we compute all values of $\Pi$ analytically.
2.2 Conditional strategies and the opportunity cost of verifying a co-player’s actions
In repeated games, agents can condition their strategies on the previous actions of their co-players. Here, we introduce an opportunity cost, denoted by $\epsilon$, which an agent must pay to know whether their partner has just made a cooperative action. The standard Tit-for-Tat (TFT) strategy (“Cooperate on the first round; in subsequent rounds, do the action that the co-player did on the previous round”) pays this cost every round. At the other extreme, fully unconditional strategies, i.e. always cooperate (ALLC) or always defect (ALLD), never pay this cost.
2.2.1 Trust-based strategies
In previous work, we introduced the concept of trust-based strategies [24]. The TUC (Trust-based Cooperation) strategy starts out playing TFT and verifying the action of its partner every round (paying the cost $\epsilon$ to do so). However, it considers the difference between the number of cooperate versus defect moves its partner has been found to make. Once this exceeds a threshold $\theta$ (known as the trust threshold), then on each subsequent round it cooperates unconditionally with probability 1p1-p (and hence does not pay $\epsilon$). On the other hand, with probability $p$ it verifies the partner’s action on the current round. If as a result the partner is found to defect, it reverts to playing TFT for the remaining rounds (see also Figure 1 of [24]).
We also consider a strategy TUD (trust-based defection), which is designed to exploit TUC. The TUD strategy begins by cooperating, but once the trust threshold $\theta$ is reached, switches to playing defect unconditionally (see also Figure 2 of [24])
The payoff matrix for the five strategies ALLC, ALLD, TFT, TUC and TUD can be written as follows:
$\bordermatrix{~&\textbf{AllC}&\textbf{AllD}&\textbf{TFT}&\textbf{TUC}&\textbf{TUD}\cr\textbf{AllC}&1&S&1&1&\frac{\theta+(r-\theta)S}{r}\cr\textbf{AllD}&T&0&\frac{T}{r}&\frac{T}{r}&\frac{T}{r}\cr\textbf{TFT}&1-\varepsilon&\frac{S}{r}-\varepsilon&1-\varepsilon&1-\varepsilon&\frac{\theta+S}{r}-\varepsilon\cr\textbf{TUC}&1-\frac{\theta\varepsilon}{r}-\frac{p(r-\theta)\varepsilon}{r}&\frac{S}{r}-\varepsilon&1-\frac{\theta\varepsilon}{r}-\frac{p(r-\theta)\varepsilon}{r}&1-\frac{\theta\varepsilon}{r}-\frac{p(r-\theta)\varepsilon}{r}&\Pi_{TUC,TUD}\cr\textbf{TUD}&\frac{\theta+(r-\theta)T-\theta\varepsilon}{r}&\frac{S}{r}-\varepsilon&\frac{\theta+T-\theta\varepsilon}{r}&\Pi_{TUD,TUC}&\frac{\theta-\theta\varepsilon}{r}\cr}.$ (2)
For ease of exposition, in the above payoff matrix the payoff for TUC interacting with TUD is written as $\Pi_{TUC,TUD}$, and the payoff for TUD against is written as $\Pi_{TUD,TUC}$. These are defined as follows (as in [24]),
$\Pi_{TUC,TUD}$ $=\frac{\theta(1-\varepsilon)}{r}+\frac{1}{rp}\left[S\left(1-(1-p)^{\,r-\theta}\right)-\varepsilon\left((1-p)^{\,r-\theta}+(r-\theta)p-1\right)\right],$
$\Pi_{TUD,TUC}$ $=\frac{1}{r}\left[\theta(1-\varepsilon)+\frac{T\left(1-(1-p)^{r-\theta}\right)}{p}\right].$
These can be understood as follows. In the first $\theta$ rounds, before the trust threshold is reached, both TUC and TUD play C and keep verifying their partners’ actions. In each of these rounds they therefore both obtain a payoff $R-\epsilon$. Once the trust threshold, $\theta$, is reached, TUC will verify the action of its partner only occasionally with probability $p$, while TUD will switch to defecting. Suppose that in the next round TUC verifies its partner’s action, and so finds it to be defecting. In this case, it obtains $S$ in the current round and $P-\epsilon$ in the remaining rounds, since it reverts to TFT. Otherwise, i.e. if TUC does not verify in that round (with probability 1p1-p), the process above is iterated for the payoffs calculation.
2.2.2 Error rate
We also allow for the possibility that agents can make unintentional errors in their actions. We assume that with probability $\mu_{\mathrm{e}}$, an agent performs the action opposite to the one given by its strategy (i.e. defects instead of cooperates, or vice versa). Figure 1 illustrates how both TUC and TFT respond to these errors. As the figure illustrates, TUC can be more resilient to these errors if they happen after the trust threshold has been reached, since they will be ignored if they occur on rounds where TUC does not verify.
Figure 1: Trust-based Cooperation (TUC) can be more robust to errors than Tit-for-Tat (TFT). The top panel shows a TUC agent playing against a TFT agent. Both agents start cooperating, and the trust threshold is reached by TUC. After a period of time, TFT makes an error and defects by mistake. However, because TUC has not verified its co-player’s action on this round, this error is overlooked and mutual cooperation resumes for a number of rounds. At a later stage, TFT defects by mistake again. This time, TUC does verify, and reverts to playing TFT for the remaining rounds. By contrast, the lower pane shows that when TFT plays against itself, the first error always causes the breakdown of mutual cooperation.
2.3 Evolutionary dynamics in finite populations
In this section, we follow the presentation of the method in Section 2.4 of [24]. To determine conditions under which trust-based strategies are advantageous, we use EGT [29, 28]. Unlike classic game theory, EGT avoids the need to assume that agents are fully rational, but instead considers a population of agents that interact (play the repeated game), receive a total payoff (also called fitness), and can then change their strategy through payoff-biased social learning. This gives rise to evolutionary dynamics where more successful strategies, with higher average payoffs from the game, increase in frequency over time as a result of being copied by less successful agents.
We model this social learning process using a pairwise comparison rule in which an agent $A$ (with fitness $f_{A}$), copies the strategy of another agent $B$ (with fitness $f_{B}$) with a probability given by the Fermi function [30],
$P_{A\rightarrow B}=\left(1+e^{-\beta(f_{B}-f_{A})}\right)^{-1}.$
The parameter $\beta$ gives the intensity of selection, i.e. how strongly an agent bases its decision to copy the strategy of another agent on the fitness difference between itself and the other agent. When $\beta=0$ the decision to copy the other agent’s strategy is random, but as $\beta$ increases copying a higher fitness strategy becomes more deterministic.
In this paper, we consider evolution in a finite population of $N$ agents. In finite populations, stochastic effects can have significant effects on evolutionary dynamics [31, 32, 33, 34]. Because of stochasticity, lower-payoff strategies may sometimes spread through the population by chance, whereas higher-payoff strategies may die out. This stochastic framework is effective in accounting for empirical findings from human behavioural experiments [32, 33], making it an appropriate method to analyse our model.
Without behavioural exploration or mutations, the end states of evolution must be monomorphic, with only one strategy present. Whenever such a state is reached, it cannot be escaped via social learning. Therefore, we also assume that, with some mutation probability, an agent can freely explore its behavioural space and hence switch to a strategy not currently being used by other individuals in the population. If we take the limit of small mutation rates then the evolutionary dynamics can be described by a Markov Chain, where each state represents a monomorphic population, and the transition probabilities between states are given by the fixation probability of a single mutant. This Markov Chain has a stationary distribution, which gives the average time the population spends in each one of these monomorphic end states. Notably, this approximation remains valid even beyond the strict limit of infinitely small mutation or exploration rates [30, 34, 35].
Let us suppose that there are at most two strategies present in the population, e.g. $k$ agents using strategy A ($0\leq k\leq N$) and (Nk)(N-k) agents using strategy B. We denote by $\pi_{X,Y}$ the payoff an agent using strategy $X$ obtains in an interaction with another agent using strategy $Y$ (as defined by the payoff matrix (2)). The average payoff to an agent that uses strategy A (B, respectively) can be written
$\begin{split}\Pi_{A}(k)=\frac{(k-1)\pi_{A,A}+(N-k)\pi_{A,B}}{N-1},\\ \Pi_{B}(k)=\frac{k\pi_{B,A}+(N-k-1)\pi_{B,B}}{N-1}.\end{split}$ (3)
The probability of changing the number $k$ of agents using strategy A by ± $\pm\ 1$ in each time step as a result of social learning is given by [30]
$T^{\pm}(k)=\frac{N-k}{N}\frac{k}{N}\left[1+e^{\mp\beta[\Pi_{A}(k)-\Pi_{B}(k)]}\right]^{-1}.$ (4)
The fixation probability of a single strategy A mutant in a population of (N1)(N-1) strategy B agents is [30, 36, 37]
$\rho_{B,A}=\left(1+\sum_{i=1}^{N-1}\prod_{j=1}^{i}\frac{T^{-}(j)}{T^{+}(j)}\right)^{-1}.$ (5)
For a set $\{1,...,s\}$ of distinct strategies, these fixation probabilities determine the transition matrix of the Markov Chain: $M=\{T_{ij}\}_{i,j=1}^{s}$, with $T_{ij,j\neq i}=\rho_{ji}/(s-1)$ and $T_{ii}=1-\sum^{s}_{j=1,j\neq i}T_{ij}$. The normalised eigenvector of the transpose of $M$ associated with the eigenvalue 1 gives the stationary distribution described above [37]. This defines the relative time the population spends adopting each strategy.
3 Results
We begin our analysis by clarifying the role of reciprocal strategies in promoting cooperation across the range of pairwise social dilemmas defined by the Prisoner’s Dilemma, Snowdrift and Stag-Hunt games. We then determine how reciprocal strategies are affected by the realistic assumption of an opportunity cost of verifying the partner agent’s actions, $\epsilon>0$, and how the effect of the opportunity cost differs between games. From this, we can then proceed to analyse the evolutionary advantage of trust, and its effectiveness in promoting cooperation across the social dilemma games. We initially assume, unless otherwise specified, that $\mu_{\mathrm{e}}=0$, i.e. agents do not make errors in their actions, unless otherwise specified.
3.1 When does reciprocity promote cooperation in different social dilemmas?
Figure 2: Frequency of the Tit-for-Tat (TFT) strategy (left), and the impact on the frequency of cooperation of adding TFT to the pool of available strategies AllC and AllD(right). The results are shown as a function of temptation to defect $T$ and sucker’s payoff $S$, and in the absence ($\epsilon=0$, top panels) or presence ($\epsilon=0.25$, bottom panels) of an opportunity cost of verifying the partner’s actions. Prisoner’s Dilemma (PD) games occur in the lower right quadrant of $S$-$T$ space, Snowdrift (SD) games occur in the top right quadrant and Stag-Hunt (SH) games in the lower left quadrant. The red dots show the particular game instances that we use for later analysis. Parameters: a game with $r=50$, and evolutionary dynamics with $\beta=0.1$ and $N=100$.
3.1.1 The role of reciprocity across social dilemma games
We first examine how Tit-for-Tat (TFT) performs across the space of Prisoner’s Dilemma (PD), Snowdrift (SD), and Stag-Hunt (SH) games, and compute how much introducing TFT into a strategy pool that initially consists only of Always Cooperate (AllC) and Always Defect (AllD) promotes cooperation. We start by considering the case where there are no opportunity costs to verifying the partner agent’s actions. Our results, presented in the top part of Figure 2, show that TFT is a frequent strategy across all the social dilemma games, and is more frequent when the temptation to defect, $T$, is high. TFT is at the highest frequency in PD types of interaction, and less frequent in SD and in SH situations. The top right panel in Figure 2, indicates that adding TFT into the pool of available strategies also increases the frequency of cooperation across all three social dilemma games.
3.1.2 Tit-for-Tat and the opportunity cost of verifying a partner’s actions
Conducting the same analysis, but this time in the presence of an opportunity cost ($\epsilon=0.25$), our results (bottom panels of Figure 2) show that including the opportunity cost drastically reduces the frequency of TFT (compare the top left and bottom left panels). The drop in frequency is the strongest in SH situations, where TFT is close to absent for many SH games in $S$-$T$ space. By contrast, the drop in frequency of TFT is smallest for PD games. Opportunity costs also reduce the effect of the presence of TFT on increasing cooperation across all games. This effect is less pronounced in SH games, even though the frequency of TFT is lowest in this part of $S$-$T$ space compared to the other games. Thus, even a low frequency of TFT promotes cooperation in the risky coordination situations represented by the SH game.
3.1.3 Cooperation results from Tit-for-Tat in Prisoner’s Dilemma games, and from Always Cooperate and Tit-for-Tat in Stag-Hunt and Snowdrift games
To understand these findings, we look more closely at differences in the role played by reciprocity in these different games. Because cooperation results from either TFT or AllC individuals cooperating with one another, TFT strategies can promote cooperation in two ways. First, TFT can directly promote cooperation by simply being a frequent strategy. Second, it can indirectly promote cooperation by catalysing the evolution of the unconditional AllC strategy.
Figure 3 examines which strategies foster cooperation in different games, and demonstrates that in the PD, cooperation primarily arises from the high frequency of TFT. In contrast, in the SH and SD games, cooperation is due to a mix of AllC and TFT strategies. When considering opportunity costs of verifying the partner agent’s actions, this distinction becomes even more pronounced, with nearly all cooperation in the SH game resulting from the higher frequency of AllC. This finding explains why opportunity costs have high impacts in the PD, where cooperation results directly from TFT individuals cooperating with each other. But there is less impact in the SH, where cooperation results from AllC players that are immune to these costs since their behaviour is unconditional on their partner’s actions. Now that we have established how TFT fares across different social dilemmas, and how a cost of verifying partner’s action affects this, we now conduct the same analysis for trust-based strategies.
Figure 3: Frequency of cooperators using the Tit-for-Tat strategy, in the absence of the opportunity cost of verifying the partner’s actions (top), and with the cost $\epsilon=0.25$ (bottom). Parameters: a game with $r=50$, and evolutionary dynamics where $\beta=0.1$ and $N=100$.
3.2 Trust across games
We now investigate, as a function of the opportunity cost, (i) the frequency of the Trust-based Cooperation strategy TUC when all five strategies are possible, and (ii) the impact of adding trust-based strategies (TUC and TUD) to the pool of available strategies on the frequency of cooperation.
Figure 4: Left: Frequency of strategies as a function of the opportunity cost, $\epsilon$. Right: Frequency of cooperation in the absence or presence of trust-based strategies TUC and TUD, as a function of the opportunity cost, $\epsilon$. The difference in frequency of cooperation when trust-based strategies are included is shaded in green where positive and red where negative. Results are presented for one example of each of the Prisoner’s Dilemma, Snowdrift and Stag-Hunt games. Parameters: trust-based strategies with $\theta=3$ and $p=0.25$, games with $r=50$, and evolutionary dynamics where $\beta=0.1$ and $N=100$.
Figure 5: Frequency of the TUC strategy (left) and the impact of considering trust strategies on the frequency of cooperation (right), as a function of the temptation to defect $T$ and the sucker’s payoff $S$. The difference in frequency of cooperation is calculated between a system with all five strategies and a system with only AllC, AllD and TFT. The red dots correspond to the values of S and T used in Figure 4. Parameters: trust-based strategies with $\theta=3$ and $p=0.25$, games with $\epsilon=0.25$ and $r=50$, and evolutionary dynamics with $\beta=0.1$ and $N=100$.
3.2.1 Baseline analysis: Trust is a successful strategy and promotes cooperation in the repeated Prisoner’s Dilemma
In the top part of Figure 4, which focuses on the PD, our analysis replicates previous findings [24]. First, increasing opportunity costs diminish the frequency of the TFT strategy (top left panel), since TFT must pay the cost of verifying every round. Second, higher opportunity costs favour TUC compared to TFT, with TUC being the most successful strategy overall for intermediate opportunity costs in the PD. Third, the inclusion of trust-based strategies (TUC and TUD) in a system that previously included only AllC, AllD and TFT leads to an overall increase in the frequency of cooperation (top right panel). When opportunity costs are high, the AllD strategy becomes the most prevalent, but (i) TUC remains among the most common across a wide range of opportunity costs, and (ii) the inclusion of trust-based strategies always (for the studied range of opportunity costs) increases the frequency of cooperation.
3.2.2 Trust is a frequent strategy across different games
Our setup now allows us to examine if these findings generalise across different types of games, and thus different types of social interaction. Our results in Figure 4 and the left panel of Figure 5 demonstrate that Trust-based Cooperation (TUC) consistently remains a frequent strategy across the three different games, and for a broad range of opportunity costs. In other words, trust is expected to be a widespread heuristic across various types of interaction, not just in PD situations. A notable difference between the games is that trust-based strategies are highly prevalent in the PD, while AllC is almost absent. In this situation, trust is a successful strategy, but naive unconditional cooperation is not. For the SD and SH, both TUC and AllC strategies are frequent, indicating that a mix of careful trusters and naive unconditional cooperators could be observed in these more relaxed social dilemmas.
3.2.3 Trust increases the frequency of cooperation when the temptation to defect is high
Our results reveal that the impact of incorporating a trust-based strategy on cooperation differs across the various games. Specifically, while trust is always beneficial in the PD, it has a limited effect on the frequency of cooperation in SD games and in SH games. As shown in Figure 5, trust-based strategies generally increase cooperation when the temptation to defect (T) is high. However, trust does not seem to increase cooperation in SH coordination games, even though it is present at a significant frequency. This is because in coordination games, we have shown earlier that reciprocal strategies are only a springboard for individuals that always cooperate to get in place. Once in place, these unconditional strategies can enjoy a high payoff without ever paying opportunity costs.
3.3 Trust when individuals make errors
Finally, we consider cases where agents can make unintentional errors in their actions. Figure 6 shows the frequency of the different strategies and of cooperation as a function of the error rate, $\mu_{\mathrm{e}}$. Our results show that an increasing error rate decreases the frequency of TUC across games, in favour of the strategy AllD. However, our results also show that the frequency of TUC remains considerable even for quite large error rates such as 5%.
Figure 6: Frequency of strategies (left) and frequency of cooperation in the absence or presence of trust-based strategies TUC and TUD (right), as a function of the error rate, $\mu_{\mathrm{e}}$. For clarity, the difference in frequency of cooperation is shaded in green when positive and red when negative. Parameters: trust-based strategies with $\theta=3$ and $p=0.25$, games with $\epsilon=0.25$ and $r=50$, and evolutionary dynamics with $\beta=0.1$ and $N=100$.
The right hand side of Figure 6 shows that including error rates up to 10% does not lessen the beneficial effect of trust on the frequency of cooperation in all of the games, and can even increase it. For instance, TUC increases cooperation the most in the PD when the error rate is around 2%-3%. Another example is that trust-based strategies improve cooperation when the error rate is above 0 in the SH, in contrast to the results where agents do not make errors.
In other words, the presence of errors, if not too frequent, can still select for trust-based strategies and make them even more beneficial for cooperation. Figure 7 shows that these conclusions hold for games repeated for a larger number of rounds ($r=100$), and where trust-based strategies verify less frequently once the trust threshold has been reached (high trustfulness, $p=0.04$).
The effect of high trustfulness on cooperation varies with the type of game. Specifically, lowering $p$ from 0.25 to 0.04 increases cooperation in SD and SH games, but it decreases cooperation in the PD (right-hand pane of Figure 7). Thus, the optimum level of trustfulness for promoting cooperation depends on the type of social dilemma, with more relaxed dilemmas favouring greater trustfulness.
Figure 7: Difference in the frequency of cooperation when trust-based strategies TUC and TUD are included, as a function of the error rate, $\mu_{\mathrm{e}}$, for three different social dilemmas and for long interactions or high trustfulness. Parameters: trust-based strategies with $\theta=3$, games with $\epsilon=0.25$, and evolutionary dynamics with $\beta=0.1$ and $N=100$.
4 Discussion
Our results can be summarised in three points. First, we have shown that trust, as a heuristic to avoid the opportunity cost of verifying its partner’s action, is a successful strategy that can evolve in a wide range of interactions. Importantly, this happens even in the presence of exploitative strategies (TUD) tailored to take advantage of this trust.
Second, by disentangling behavioural measures of trust from behavioural measures of cooperation, we have been able to rigorously examine the effect that trust has on the overall level of cooperation in a population. Our results show that the presence of the Trust-based Cooperation strategy (TUC) increases the overall level of cooperation when the temptation to defect is high, but not in coordination dilemmas such as Stag-Hunt games. We have shown that this is because reciprocity in coordination games is useful to kickstart cooperation, but is not necessary once cooperation is established. This suggests that interventions aiming to promote the success of TUC, such as reducing high opportunity costs, would have less effect on cooperation in risky coordination games compared to stricter social dilemmas.
Third, we have shown that when there is the possibility of agents making errors in their actions, then TUC remains frequent, and can even be more beneficial for cooperation compared to the case where agents never make errors. Notably, TUC does not suffer as much from errors as TFT, as TUC agents ignore errors from cooperative partners on the rounds that they do not verify. By contrast, TFT is not trustful and verifies its partner’s action every round, which means that one single mistake causes TFT to defect in retaliation. This can then cause the partner to defect in retaliation, leading to a loss of cooperation over many rounds resulting from one mistake [38]. Trust can prevent this spiral into defection.
Our model is based on the assumption that it can be costly to know whether the action of another agent was actually cooperative – even when you observe that action [24]. Everyday examples of this are abundant. For example, there is a time cost to verifying if the correct amount has been debited from your bank account, or if goods that you have ordered online are genuine rather than fake. When interacting with artificial intelligence (AI), these costs can become even greater. For example, are personalised music recommendations from a streaming app influenced by paid promotions [39]? Has an LLM agent completed a task in a way that meets the preferences I prompted it with, or were its actions not in my interests due to biases in its training data [40]?
Our operationalisation of trust – as choosing not to verify another agent’s actions – provides an objective behavioural measure that applies across a broad spectrum of social interactions. Using this, trust can be measured in experimental settings by how often someone chooses (or not) to pay a cost to verify another agent’s actions. This could be carried out in standard behavioural economics settings [41], where a participant is given the option to pay to find out the action of its partner in each round of a repeated social dilemma game. Some experimental studies have already considered costly verification for the particular case of the “trust game” [42, 43]. In this game, one player is the investor and the other player is the trustee. The investor begins with an endowment amount of money. They then choose a proportion (between 0 and 1 inclusive) of this to send to the trustee. The amount sent is multiplied by a factor >1>1. The trustee then chooses the proportion of this to return to the investor. When the game is repeated, most experiments assume that the investor is given complete information about the previous actions of the trustee at no cost. However, a study has shown that when a cost to obtain this information is introduced, people are sensitive to this cost and verify less [42]. Moreover, the frequency of verifying was lower in later rounds, in a pattern akin to the TUC strategy. Future work should extend these experiments beyond the trust game and into the larger space of social dilemma games that we have modelled.
Crucially, this trust measure can also be applied in more realistic experimental settings. One important application of this is experiments to measure the factors affecting people’s trust in AI. Although there have been many empirical studies in computer science that have aimed to measure factors affecting trust in various applications of AI, these studies have lacked a unified theoretical framework [44, 45]. This has led to conflicting results and a lack of predictive models, with “Trustworthiness mechanisms and measures […] being advanced in AI regulations and standards that may not actually increase trust” [46]. The trust measure we have introduced here can help to resolve this impasse, by providing an objective measure of trust that is grounded in (evolutionary) game theory. For example, incentivised experiments could be designed where a user can take advice from an AI agent in a realistic setting, such as an e-commerce chatbot recommending products or an LLM answering questions on a text, and can pay a cost to verify whether that advice is correct. This can then be connected with EGT models by varying factors such as how frequently the AI gives correct information (i.e. its strategy to cooperate or defect), or whether the AI has a high reputation from other users.
The costs of verifying also lead to problems for the alignment and control of AI, that is, ensuring that the goals and behaviours of AI systems are aligned with human values and intentions. Research in AI safety has stressed the risk of scheming – AI systems that deceptively appear to be aligned with human goals during training, but then pursue misaligned goals after deployment once they are no longer closely monitored [47]. This type of scheming corresponds to the TUD strategy in our model – cooperate until an auditor or developer relaxes oversight, and then defect. The opportunity cost of verifying maps to the cost of auditing the AI system. Extensions of our model, with the payoff matrix calibrated to the parameters used in red-team blue-team AI control experiments [48], could therefore provide insights into auditing regimes and other institutional incentives that minimise vulnerability to scheming by LLMs and other future AI systems.
Most generally, our model demonstrates the evolutionary success of a trust heuristic in competition with strategies that do not use trust (ALLC, ALLD and TFT), and in competition with strategies designed to build and then exploit trust (TUD). Trust evolves even though trust-based strategies are costly to use compared to unconditional strategies (since they pay the cost of verifying before trust is established and periodically thereafter), and are risky compared to standard reciprocal strategies (since they could potentially be exploited once they stop verifying every round). These results suggest that a trust heuristic could have evolved in humans by natural selection. Indeed, there is evidence that human trust behaviour is to some extent genetically determined [49], and hence is an evolved trait. Our results suggest that repeated social dilemma games, as played by humans throughout (pre)history [50], could provide selection pressure for the evolution of trust.
There are several limits to our model which could be addressed in future work. First, we have not modelled the evolution of the parameters of trust-based strategies that control when they verify their co-player’s actions, $\theta$ and $p$. Future work could allow these to be evolving traits, allowing evolutionary optimal values to be found for different game parameters.
Second, we assumed that lost trust cannot be gained again. In the real world, trust relationships can be rebuilt if interactions become cooperative again. This could potentially allow for greater long-term cooperation than is seen in our results, but could also open the door for more exploitative TUD style strategies. Exploitative strategies that take advantage of trust rebuilding would not necessarily be successful, though, because they need to cooperate for a long period in order to rebuild trust before being able to exploit again. Moreover, this dynamic could drive the evolution of the parameters of Trust-based Cooperation, for example requiring a longer period of cooperation before trust is built (greater $\theta$), and a lower trustfulness (higher $p$). An important next step is therefore to consider the effect of more complex deceiving strategies on the form of trust that emerges.
Third, we have modelled a well-mixed population of agents that are all similar in their cognitive capabilities. It would be important to examine the effects of spatial structure [51], which can facilitate the evolution of cooperation in repeated games such as the Prisoner’s Dilemma [23]. Moreover, agents might differ in how easily they can verify another agent’s actions, and in the complexity of strategies that they can use. For example, when considering interactions between humans and AI agents, the model could be extended to allow AI agents to verify actions at lower costs, and to use more complex strategies such as modelling and predicting human-partner trust behaviours. This would represent AI agents’ greater access to information and processing power, respectively. This could be achieved by modelling a separate population of AI agents that compete with and copy strategies from each other, but earn payoffs from playing with a population of human agents.
Finally, the model of trust as a cognitive shortcut to avoid the cost of verification could apply beyond interactions that are repeated between the same pair of agents. These include $n$-player social dilemmas such as repeated public goods games. Trust-based strategies could also apply with indirect reciprocity, in cases where it is costly to get information from third parties about a partner’s previous actions.
In conclusion, we have developed a behavioural measure of trust applies across the spectrum of symmetric two-player social dilemma games, and that distinguishes observed trust from observed cooperation. Because this behavioural measure does not depend on access to an agent’s internal state, it can apply to both human and AI agents. Using the methods of evolutionary game theory, we have shown that Trust-based Cooperation, which verifies its partners’ actions only occasionally after a trust-threshold is reached, is a successful strategy in a wide range of social interactions. We have also shown that the presence of trust increases the overall level of cooperation in a population, especially when agents can make unintentional errors in their actions. This adds formal support for the idea that high-trust societies lead to more cooperative and friction-free economies [1].
Declaration of generative AI and AI-assisted technologies
in the manuscript preparation process
The authors used an AI-based tool (ChatGPT 5.1 with Deep Research assistance) solely to support the identification of relevant literature on costly monitoring in classic game theory and behavioural economics. All cited sources were independently reviewed and integrated by the authors. No AI-generated text was used in the submitted manuscript.
Acknowledgements
T.A.H. is supported by EPSRC (grant EP/Y00857X/1).